### v1

# On the interplay between inverse scattering for asymptotically hyperbolic manifolds and the Calderón problem for the Conformal Laplacian

Abstract.
In this short note, we use the relation obtained by Guillarmou–Guillopé and Chang–González between the generalized eigenvalue problem for asymptotically hyperbolic (AH) manifolds and the Conformal Laplacian, to obtain a new inverse scattering result: on an AH manifold of dimension $n+1$, we show that the scattering matrix at energy $\frac{n+1}{2}$ determines the jet of the metric on the boundary, up to a diffeomorphism and conformal factor.
1. Introduction
Let $\overline{X}$ be a compact manifold with boundary of dimension $n+1$. A smooth metric $g$ in the interior $X$ of $\overline{X}$ is conformally compact (CC) if $\overline{g}=\rho^{2}g$ extends smoothly to a Riemannian metric in $\overline{X}$, where $\rho\in C^{\infty}(\overline{X},[0,\infty))$ is a boundary defining function for $\partial\overline{X}$ (i.e., $d\rho|_{\partial\overline{X}}\neq 0$ and $\{\rho=0\}=\partial\overline{X}$). The boundary $\partial\overline{X}$ equipped with the conformal class of $\overline{g}|_{T\partial\overline{M}}$ is called the conformal boundary or conformal infinity of $(\overline{X},g)$. If $|d\rho|_{\overline{g}}=1$ at the boundary, $g$ is called asymptotically hyperbolic (AH). A more restrictive condition is asking for the Einstein condition $\operatorname{Ric}_{g}=-ng$. In this case, $g$ is called conformally compact Einstein (CCE or Poincaré–Einstein).
In these particular “edge” geometries (in the sense of [Mazzeo91_ell]), we study the inverse scattering problem for the generalized eigenvalue problem
$-\Delta_{g}u-s(n-s)u=0\quad\text{in }X,$ (1.1)
where $s=\frac{n+1}{2}$. If $(\overline{X},g)$ is an AH manifold with $g$ even modulo $O(\rho^{3})$, given $f\in C^{\infty}(\partial\overline{X})$ the problem (1.1) has a unique solution $u$ with $u|_{\partial\overline{X}}=f$ and the expansion
$u=\rho^{n-s}(f+O(\rho^{3}))+\rho^{s}\left(-S\left(\frac{n+1}{2}\right)+O(\rho)\right),$
see [GZ03, Guillarmou05, GG07] and Lemma 2.2 below. $S$ (sometimes written as $S_{g}$ to make clear the dependence on $g$) is the scattering matrix at energy $\frac{n+1}{2}$, and depends on the choice of the boundary defining function (or equivalently, of the choice of the metric on the conformal infinity), although it can be defined without this dependence, see [GZ03]*Section 3). Hereafter, we denote by $P_{s}$ the operator appearing in (1.1).
As is mentioned in [MM87, GZ03], the scattering matrix can be considered as a version of the Dirichlet-to-Neumann (DN) for these geometries. Then, it then makes then sense to consider the analog of the Calderón problem to this context. In inverse problems, this means to determine the metric (or even the differential or topological structure) of a compact Riemannian manifold with boundary, from the DN map associated to the Laplace–Beltrami operator. On compact Riemannian manifolds, this is solved on surfaces and in the analytic category [LU01], and in conformally anisotropic geometries [DSFKSU09]. The Calderón problem in our context, known as the inverse scattering problem, asks if it is possible to recover information of the manifold $(\overline{X},g)$ from the knowledge of the scattering matrix $S_{g}(s)$, for certain fixed $s$. In this direction, our main result is the following:
Theorem 1.1.
Let $(\overline{X}_{i},g_{i})$ be asymptotically hyperbolic manifolds of dimension $n+1\geq 3$, where the metrics are even modulo $O(\rho^{3})$ (see Definition 2.1). Suppose that $\partial\overline{X}_{i}$ contain a common open set $\Gamma$ such that the identity map $id\colon\Gamma\subset\partial\overline{X}_{i}\to\Gamma\subset\overline{X}_{2}$ is a diffeomorphism. Assume that $\ker_{L^{2}}P_{\frac{n+1}{2}}^{i}=\emptyset$ for $i=1,2$. If for some boundary defining functions $\rho_{i}$, the scattering matrices at energy $\frac{n+1}{2}$ agree for the two metrics, i.e., $S_{g_{1}}(\frac{n+1}{2})=S_{g_{2}}(\frac{n+1}{2})$, then there exists $c\in C^{\infty}(\overline{X}_{1})$ such that:
  1. (1)
    when $n=1$, there exists a diffeomorphism $\psi\colon\overline{X}_{1}\to\overline{X}_{2}$ with $\psi|_{\Gamma}=id_{\Gamma}$ and such that $c\psi^{*}g_{2}=g_{1}$;
  2. (2)
    when $n\geq 2$, there exists a diffeomorphism $\psi\colon U_{1}\to U_{2}$ between neighborhoods $U_{i}$ of $\Gamma$ in $\overline{X}_{i}$, with $\psi|_{\Gamma}=id_{\Gamma}$ and such that $c\psi^{*}g_{2}$ and $g_{1}$ coincide up to infinite order.
The result is inspired by the classical result for compact Riemannian manifolds with boundary obtained in [LU89]. In contrast to the proof given in [LU89], which is based on a factorization of the Laplace–Beltrami operator, our proof follows an indirect approach (although we also have a “formal” factorization of the Laplacian in AH geometries, see Lemma 2.11): the relation between the Dirichlet problem for (1.1) at $s=\frac{n+1}{2}$ and the Dirichlet problem for the Conformal Laplacian (also known as the Yamabe operator), described first in [GG07] for AH manifolds and then in [CG11] for more general equations in CCE manifolds. Note that this means we have conformal invariance for the scattering operator this energy, so the conformal factors appearing in the result are natural. They show that, for manifolds as in Theorem 1.1, the DN map of $\rho^{2}g$ is the scattering of $g$ at energy $\frac{n+1}{2}$. According to the results of [LLS22], the DN map for the Conformal Laplacian determines the jet of a metric at the boundary, which is conformal to $\rho^{2}g$. We remark, however, that the determination of the jet on the boundary in [LLS22] relies on a factorization of the magnetic Schrödinger operator carried out in [DSFKSU09].
Although the theorem has a short proof, the result is new since there are no inverse scattering results at this energy level, and furthermore, we impose no restrictions on the dimension. To the authors’ knowledge, the first inverse scattering result in the AH manifold context was given in [JSB00], where it was shown that the scattering matrix on AH manifolds associated with (1.1) with a potential determines the metric up to diffeomorphism at infinite order, assuming the same potential, and also the potential up to infinite order if one fixes the metric. They results do not apply in our case due to their restrictions on the energy levels. Using radiation fields, it was proven in [SB05] that the scattering matrix at all energies determines the metric up to a diffeomorphism on AH manifolds. In the same spirit, inverse scattering results for asymptotically complex hyperbolic manifolds were obtained in [GSB08]. On even dimensional CCE manifolds (of dimension $n+1$), it is known [GSB09] that the scattering matrix at energy $n$ determines the metric up to a diffemorphism. On CC manifolds, it was shown in [Marazzi09] that the scattering matrix at two energy levels, associated with (1.1) with potential, determines the potential and the metric (and even the curvature at the boundary) up to the usual gauge. In [GT11], an inverse scattering result for the Schrödinger equation on AH manifolds was obtained, showing that the scattering matrix determines the metric, assuming that the potentials vanish up to order 2 in the boundary. Below, we mention an updated version of this result, which also recovers the metric, see Theorem 2.13. The results from [GSB09] were extended in [Marazzi11], where it is shown that on CCE manifolds, the scattering matrix at energy $n$ determines the metric at certain order involving $\log$ terms, up to diffeomorphism, and a modified scattering operator determines the metric up to diffeomorphism. In [Isozaki04], inverse scattering problems in the hyperbolic space were studied and related, similarly to this work, to inverse problems for the Schrödinger operator in Euclidean domains. In [IKL10], inverse scattering was consider for manifolds with cylindrical ends, while in [IKL17], surfaces with conical singularities, cusps and regular ends were considered, see also the monographs [IK14, IKL14, IKL15]. The results in [SB05] were extended to partial data in [HSB15], and for scattering with disjoint source and observation sets in [HSB18]. We also highlight the work [Marx-Kuo24] on AH manifolds, where, assuming some evenness conditions on the metric, it is shown that one can determine it from the knowledge of minimal areas.
Acknowledgments
The author would like to thank Colin Guillarmou, Rafe Mazzeo, Mikko Salo, Gunther Uhlmann, and András Vasy for helpful discussions. Additionally, the author expresses gratitude to Stanford University and the University of Washington for their hospitality during his visit, while this work was in progress. The author was partly supported by the Ross-Lynn Scholar Grant from Purdue University’s College of Science.
2. Proof of the Main Result
Before beginning the analysis, we recall the definition of evenness in the sense of Guillarmou.
Definition 2.1.
An asymptotically hyperbolic manifold $(M,g)$ is even modulo $O(\rho^{2k+1})$ if there is a boundary defining function $\rho$ with $|d\rho|_{\rho^{2}g}=1$, and such that in a collar neighborhood, the pullback of $g$ by the gradient flow of $\nabla^{\rho^{2}g}\rho$ takes the form
$\frac{1}{t^{2}}\left(dt^{2}+\sum_{i=0}^{k}h_{2i}t^{2i}+h_{2k+1}t^{2k+1}+O(t^{2k+2})\right),$
where $h_{j}$’s are symmetric tensors on $\partial\overline{X}$.
It was shown in [Guillarmou05]*Lemma 2.1 that this is well-defined, meaning that this notion is independent of the boundary defining function. Examples of even metrics can be found in the remark after Theorem 1.4 in [Guillarmou05] and the references therein. For instance, hyperbolic metrics perturbed on a compact set are even modulo $O(\rho^{\infty})$, and CCE manifolds of dimension $n+1$ are even modulo $O(\rho^{n})$.
Now we turn out our attention to the behavior of solutions to (1.1) near the boundary. The most fundamental work in this direction is [GZ03]. The following is mentioned in [GG07] as a consequence of [Guillarmou05]*Lemma 4.1, but we provide a short proof by completeness.
Lemma 2.2.
Let $(\overline{X},g)$ be an asymptotically hyperbolic manifold with $g$ even modulo $O(\rho^{3})$. Assume that $\ker_{L^{2}}P_{\frac{n+1}{2}}=\emptyset$. Given $f\in C^{\infty}(\partial\overline{X})$, there exists a unique solution $u$ to
$P_{\frac{n+1}{2}}u=0,\quad u|_{\partial\overline{X}}=f,$
having the expansion
$u=\rho^{\frac{n-1}{2}}(f+O(\rho^{3}))+\rho^{\frac{n+1}{2}}\left(-S\left(\frac{n+1}{2}\right)+O(\rho)\right).$
The term of order $\frac{n+1}{2}$, i.e., $S((n+1)/2)$ is defined as the scattering matrix of energy $(n+1)/2$.
Proof.
We follow [GZ03, Guillarmou05, GSB09]. The strategy is as follows: First, we construct a solution modulo infinite order. Then, the hypothesis on the kernel and the metric allow us to apply [Guillarmou05]*Lemma 4.1 to ensure that the resolvent has no poles at $(n+1)/2$. Therefore, we can modify the just constructed solution to obtain a full solution having the required asymptotics.
The first step is to construct $F_{\infty}\in C^{\infty}(\overline{X})$ such that
$P_{\frac{n+1}{2}}\rho^{\frac{n-1}{2}}F_{\infty}=O(\rho^{\infty}),\quad F_{\infty}|_{\rho=0}=f.$ (2.1)
where $\rho$ is a Graham–Lee geodesic boundary defining function (see [GL91]*Lemma 5.2 and the subsequent paragraph, also [Graham00]*Lemma 2.1). As mentioned in [Guillarmou05]*p. 24, it is enough to work in a neighborhood of the boundary. Using the flow $\phi_{t}(y)$ of the gradient $\nabla^{\rho^{2}g}\rho$ of the Graham–Lee boundary defining function, we obtain a diffeomorphism $\phi\colon[0,\varepsilon)_{t}\times\partial\overline{X}\to\phi([0,\varepsilon)\times\partial\overline{X})\subset\overline{X}$ defined by $\phi(t,y):=\phi_{t}(y)$, and the metric pulls back to
$\phi^{*}g=\frac{dt^{2}+h(t)}{t^{2}},$
for some smooth one-parameter family of metrics $h(t)$ on the boundary $\partial\overline{X}$, where $\phi^{*}\rho=t$. It can be shown that
$P_{\frac{n+1}{2}}t^{\frac{n-1}{2}}=t^{\frac{n-1}{2}}\mathscr{D},$
where
$\mathscr{D}:=-t^{2}\partial_{x}^{2}-\frac{t}{2}\operatorname{tr}_{h}\partial_{t}ht\partial_{t}-\frac{n-1}{4}\operatorname{tr}_{h}\partial_{t}h+t^{2}\Delta_{h}.$
Observe that for $j\in\mathbb{N}_{0}=\{0,1,\ldots\}$ we have
$\mathscr{D}(ft^{j})=j(1-j)ft^{j}+t^{j}G\left(\frac{n+1}{2}-j\right)f,$ (2.2)
where
$(G(z)f)(x,y):=t^{2}\Delta_{h}f-\frac{n-z}{2}t\operatorname{tr}_{h}\partial_{t}hf.$
Suppose now that $\tilde{F}\in C^{\infty}(\overline{X})$ satisfies $\mathscr{D}\tilde{F}=O(t^{j})$ for $j\geq 1$. Since $G(z)f=O(t)$ (a priori, we analyze this further below), using (2.2) with $f=(t^{-j}\mathscr{D}\tilde{F})|_{t=0}/(j(1-j))$, we obtain
$\mathscr{D}\tilde{F}-\mathscr{D}\left(t^{j}\frac{(t^{-j}\mathscr{D}\tilde{F})|_{t=0}}{j(1-j)}\right)=O(t^{j+1}),$ (2.3)
for $j\neq 1$. Indeed, on the left hand side of (2.3), we are subtracting the term of order $O(t^{j})$ from $\mathscr{D}\tilde{F}$. This procedure suggests the following recursive definition of $f_{j}$ and $F_{j}$:
$F_{0}:=f_{0},\quad f_{j}:=\frac{-(t^{-j}\mathscr{D}(F_{j-1}))|_{t=0}}{j(1-j)},\quad F_{j}:=F_{j-1}+f_{j}t^{j},\quad j\geq 4.$ (2.4)
To construct $f_{1}$, $f_{2}$, and $f_{3}$, we proceed as follows: To construct $f_{1}$, and hence $F_{1}$, we look up for terms or order $1$ in $\mathscr{D}F_{0}$. First, recall from [Guillarmou05]*Equations (4.8), (4.9) that since $\operatorname{tr}_{h_{0}}h_{1}=0$, we have
$G(z)=t^{2}(\Delta_{h_{0}}+D^{0})+t^{3}D^{2}+O(t^{4}),$
where $D^{j}$ are differential operators of order $j$ on $\partial\overline{X}$. So $\mathscr{D}f$ only contains terms of order $3$ or more. Hence, $f_{1}:=0$ and $F_{1}:=F_{0}$. Similarly, for $f_{2}$ we look up for terms or order $2$ in $\mathscr{D}F_{1}$. By the previous argument, there are non of them, and we set $f_{2}:=0$ and $F_{2}:=F_{0}$. Finally, for $f_{3}$ we see that $t^{3}(\Delta_{h_{0}}+D^{2})f$ is the term of order 3 in $\mathscr{D}F_{2}=\mathscr{D}f$, and hence we define $f_{3}:=(\Delta_{h_{0}}+D^{2})f$ and $F_{3}=t^{3}f_{3}+f_{0}$. We can now continue using (2.4) inductively. By Borel’s lemma, we can find $F_{\infty}\in C^{\infty}(\overline{X})$ such that $F_{\infty}-F_{j}=O(t^{j+1})$, and satisfying (2.1).
To complete the construction of the full solution to the eigenvalue problem, we now use our assumption on $g$ and the $L^{2}$ kernel to invoke [Guillarmou05]*Lemma 4.1, which shows that the resolvent of $P_{s}$ has no poles at $s=(n+1)/2$. Hence, it is an analytic family by [GZ03]*Proposition 3.1 (and the results in [MM87]), the resolvent of $P_{\frac{n+1}{2}}$ satisfies
$R\left(\frac{n+1}{2}\right)\colon\dot{C}^{\infty}(\overline{X})\to x^{\frac{n+1}{2}}C^{\infty}(\overline{X}),$
where $\dot{C}^{\infty}(\overline{X}):=\{u\in C^{\infty}(\overline{X}):u=O(x^{\infty})\}$. Therefore, we can set
$u=x^{\frac{n-1}{2}}F_{\infty}-R\left(\frac{n+1}{2}\right)x^{\frac{n-1}{2}}F_{\infty}$
which solves the eigenvalue problem.
To see the uniqueness, we notice that if $u_{1}$ and $u_{2}$ are two solutions with the same term of order $\rho^{\frac{n-1}{2}}$, then $u_{1}-u_{2}\in\rho^{\frac{n+1}{2}}C^{\infty}(\overline{X})$. This implies that $u_{1}-u_{2}\in L^{2}(\overline{X})$, contradicting our hypothesis. This finishes the proof. ∎
Remark 2.3.
Alternatively, one can use a more direct method in the following sense. One can obtain a solution to the problem $P_{s}u^{\prime}_{s}=O(\rho^{\infty})$ with $u^{\prime}_{s}|_{\partial\overline{X}}=f$ for $s$ close to $(n+1)/2$. Then, a full solution can be obtained by using the resolvent at energy $s$. Since there are no poles, $R(s)$ extends to $R((n+1)/2)$, and one then obtain a solution $u$ to $P_{\frac{n+1}{2}}u=0$ and $u_{|\partial\overline{X}}=f$.
Remark 2.4.
We note that we only used the vanishing of $\operatorname{tr}_{h_{0}}h_{1}$, not the full evenness condition. The vanishing of the trace means that $\partial\overline{X}$ minimal.
The key relation between inverse scattering on AH manifolds and the Calderón problem for the Conformal Laplacian is given by the following result of Guillarmou and Guilollopé:
Lemma 2.5 ([GG07]*Section 4).
Let $(\overline{X},g)$ be an asymptotically hyperbolic manifold, even modulo $O(\rho^{3})$. Then $u\in C^{\infty}(\overline{X})$ is a solution to
$P_{\frac{n+1}{2}}u=0,\quad u|_{\partial\overline{X}}=f,$
if and only if $U:=\rho^{-\frac{n-1}{2}}u$ is a solution to
$\left(-\Delta_{\overline{g}}+\frac{n-1}{4n}R_{\overline{g}}\right)U=0,\quad U|_{\partial\overline{X}}=f.$
Furthermore,
$S\left(\frac{n+1}{2}\right)=\Lambda_{\overline{g}},$
as operators from $C^{\infty}(\partial\overline{X})$ to itself.
We point out that on CCE manifolds, this was extended to any energy level in [CG11]*Lemma 4.1, Theorem 4.3.
Remark 2.6.
We emphasize that in [GG07], this is mentioned assuming that the scalar curvature of $(M,g)$ is constant equal to $-n(n+1)$. However, this is used to obtain $\operatorname{tr}_{h_{0}}h_{1}=0$ to make this result work, which, as we saw before, is implied by the metric being even modulo $O(\rho^{3})$. Thus, in particular, the above result is true for any CCE manifold, since they have constant scalar curvature $-n(n+1)$, which implies $\operatorname{tr}_{h_{0}}h_{1}=0$, see [GG07]*Section 4. Therefore, lemmas 2.2 and 2.5 apply to any CCE manifold.
To reach our goal, we will need the following result concerning the boundary jet recovery in the inverse problem for the Conformal Laplacian.
Lemma 2.7 ([LLS22]*Lemmas 2.1 & 2.2).
Let $(M_{1},\overline{g}_{1})$, $(M_{1},\overline{g}_{2})$ be compact Riemannian manifolds with boundary, with a common boundary portion $\Gamma\subset M_{1},M_{2}$. Assume that Dirichlet-to-Neumann maps for the Conformal Laplacian agree, i.e., $\Lambda_{\overline{g}_{1}}=\Lambda_{\overline{g}_{2}}$. Then, there exist $c_{i}\in C^{\infty}(M_{i})$ positive functions with $c_{i}|_{\Gamma}=1$ such that if $\tilde{g}_{i}=c_{i}\overline{g}_{i}$, then the jets of $\tilde{g}_{1}$ and $\tilde{g}_{2}$ on the respective $\tilde{g}_{i}$-normal coordinates agree.
Proof of Theorem 1.1.
Let $\rho_{i}$ be geodesic boundary defining functions for $g_{i}$. Since the scattering matrices at energy $\frac{n+1}{2}$ coincide, Lemma 2.5 implies that the same holds for the DN maps associated to the Conformal Laplacian of the metrics $\rho_{1}^{2}g_{1}$ and $\rho_{2}^{2}g_{2}$. If $n\geq 2$, Lemma 2.7 says that there exist $c_{i}\in C^{\infty}(\overline{X}_{i})$ with $c_{i}|_{\Gamma}=1$, $i=1,2$ such that given coordinates $y$ in $\Gamma$, if $\psi_{1}$ and $\psi_{2}$ are $c_{1}\rho_{1}^{2}g_{1}$- and $c_{2}\rho_{2}^{2}g_{2}$-boundary normal coordinates constructed by using the same $y$-coordinates, then $(\psi_{1}^{-1})^{*}(c_{1}\rho_{1}^{2}g_{1})=(\psi_{2}^{-1})^{*}(c_{2}\rho_{2}^{2}g_{2})+O(t^{\infty})$. Hence, $g_{1}=(c_{1}\rho_{1}^{2})^{-1}(\psi_{1}\circ\psi_{2}^{-1})^{*}(c_{2}\rho_{2}^{2}g_{2})+O(\rho_{1}^{\infty})$. To deal with surfaces, one applies the result from [LU01] instead of [LLS22] to conclude the proof. ∎
Remark 2.8.
Note that for surfaces, the hypothesis in the kernel is superfluous by the results in [MM87, GZ03].
Remark 2.9.
By the results in [GG07], one actually obtains equivalence between the inverse scattering problem at energy $1$ and the Calderón problem for surfaces. However, in higher dimensions, one also needs information of the curvature of the boundary with respect to the compactified metric.
Remark 2.10.
When $n+1\geq 2$, if all the components are analytic, i.e., the metrics, the conformal factors and the diffeomorphisms, then one can extend the diffemorphism as in [LU89] to obtain a global one. Furthermore, if one assumes that the compactified manifolds $(\overline{X}_{i},c_{i}g_{i})$ are locally conformally real analytic in the sense of [LLS22], then a global result can be achieved as in the case of surfaces.
We would like to point out that one needs the conformal factors in order to obtain a vanishing condition in a certain expression involving the metric tensor in order to recover the jet of the metric at the boundary. The recovery of the jet is obtained by the following procedure, see [DSFKSU09]*Equation 8.2, [LLS22]*p. 11136111We point out that there is a little typo there, the condition should be $\partial_{\tilde{x}_{n}}^{j}(\tilde{g}_{\alpha\beta}\partial_{\tilde{x}_{n}}\tilde{g}^{\alpha\beta})(x^{\prime},0)=0,\quad j\geq 1,$ First, one factorizes the Conformal Laplacian, and then one shows that the Dirichlet-to-Neumann map is equal to a pseudodifferential operator of order 1 appearing in such factorization, modulo a smoothing operator. Hence, to recover the jet of the metric, one shows that the DN map determines this $\Psi$DO, and finally one shows that from the full symbol one can recover the jet of the metric. In the context of inverse problems, this technique was first used in [SU88] for the conductivity equation, [LU89] for the Laplace–Beltrami operator, and then [NSU95] for the magnetic Schrödinger operator with Euclidean metric. In [DSFKSU09] the procedure was generalized for the Schrödinger operator in the presence of a magnetic field on Riemannian manifolds. We observe that if in particular we have no magnetic field this procedure gives a (local) factorization of $P_{s}$. Indeed, in the coordinates given by the geodesic boundary defining function $\rho$ (i.e., on $\partial\overline{X}\times[0,\varepsilon)$, we use coordinates $(y,t)$ so that $\overline{g}=\rho^{2}g=dt^{2}+h_{t}(y)$, with dual variables $(\xi,\eta)$), we have
$L_{\overline{g}}=(D_{t}+iE-iB(y,t,D_{y}))(D_{t}+iB(y,t,D_{y}))$ (2.5)
modulo $\Psi^{-\infty}$,
$E(x)=\frac{1}{2}\overline{g}_{\alpha\beta}\partial_{t}\overline{g}^{\alpha\beta},$
where $B\in\Psi^{1}$ has symbol $b\sim\sum_{j\leq 1}b_{j}$, where
$b_{0}$ $=-\sqrt{q_{2}},$
$b_{1}$ $=\frac{1}{2b_{1}}(-\partial_{t}b_{1}+Eb_{1}-\nabla_{\eta}b_{1}\cdot D_{y}b_{1}+q_{1}),$
$b_{-1}$ $=\frac{1}{2b_{1}}\left(-\partial_{t}b_{0}+Eb_{0}-\sum_{\begin{subarray}{c}0\leq j,k\leq 1\\ j+k=|K|\end{subarray}}\frac{\partial_{\eta}^{K}b_{j}D_{y}^{K}b_{k}}{K!}+\frac{n-1}{4n}R_{\overline{g}}\right),$
$b_{m-1}$ $=\frac{1}{2b_{1}}\left(-\partial_{t}b_{m}+Eb_{m}-\sum_{\begin{subarray}{c}m\leq j,k\leq 1\\ j+k-|K|=m\end{subarray}}\frac{\partial_{\eta}^{K}b_{j}D_{y}^{K}b_{k}}{K!}\right)\quad(m\leq-1).$
Here, $q_{2}$ and $q_{1}$ are the symbols of
$\overline{g}^{\alpha\beta}D_{\alpha}D_{\beta},\qquad-i\left(\frac{1}{2}\overline{g}^{\alpha\beta}\partial_{\alpha}(\log|g|)+\partial_{\alpha}\overline{g}^{\alpha\beta}\right)D_{\beta},$
respectively. Here the left quantization is used. Hence, using the invariance of the Conformal Laplacian together with (2.5), we obtain
$P_{\frac{n+1}{2}}$ $=t^{\frac{n+3}{2}}L_{\overline{g}}t^{-\frac{n-1}{2}}$
$=t^{\frac{n+3}{2}}L_{\overline{g}}t^{-\frac{n-1}{2}}$
$=t^{\frac{n+3}{2}}(D_{t}+iE-iB(y,t,D_{y}))(D_{t}+iB(y,t,D_{y}))t^{-\frac{n-1}{2}}$
$=t^{\frac{n+3}{2}}(D_{t}+iE-iB(y,t,D_{y}))t^{-\frac{n+1}{2}}\left(tD_{t}-\frac{n-1}{2}+itB\right)$
$=\left(tD_{t}-\frac{n+1}{2}+itE-itB\right)\left(tD_{t}-\frac{n-1}{2}+itB\right).$
We summarize this discussion as follows:
Lemma 2.11.
Formally, we have the factorization
$P_{\frac{n+1}{2}}=\left(tD_{t}-\frac{n+1}{2}+itE-itB\right)\left(tD_{t}-\frac{n-1}{2}+itB\right)$
modulo $t\Psi^{-\infty}$.
We refer to this as a formal factorization, since it is merely an algebraic manipulation, but not a proper factorization in the standard algebra of $\Psi$DO. Since $\Delta_{g}$ is a 0-differential operator ([Mazzeo91_ell]*Proposition 2.6), one should expect to have a factorization in the 0-calculus of Mazzeo and Melrose. Lemma 2.11 suggests in general a factorization of the following form
$P_{s}=\left(tD_{t}-s+iE-iA\right)\left(tD_{t}-(n-s)+iA\right).$ (2.6)
Remark 2.12.
We point out that it is shown in [CG11] that there exists “special” boundary defining functions $\rho_{1}$, $\rho_{2}$ so that, in a collar neighborhood of the boundary, $P_{s}$ becomes:
  1. (1)
    the Laplace–Beltrami operator associated to $\rho_{1}^{2}g$, when $s=\frac{n+1}{2}$,
  2. (2)
    the Bakry–Émery Laplacian (also known as Ornstein-Uhlenbeck operator, Witten Laplacian or Weighted Laplacian) $-\operatorname{div}((\rho_{2})^{a}\nabla\bullet)$. Inverse problems for this operator have been studied recently in [BK25]. The equation can be thought as applying the operator
    $L=e^{V}\operatorname{div}(e^{-V}\nabla\bullet),$
    where $V=-\log\rho_{2}^{a}$.
Finally, we state the following stronger version of Theorem 1.1 in dimension 2.
Theorem 2.13.
Let $(\overline{X},g_{i})$ be two asymptotically hyperbolic surfaces with common boundary $\partial\overline{X}$, and let $V_{i}\in\rho_{i}^{2}C^{\infty}(\overline{X})$ be two potentials. Assume that
$\{\partial_{\rho_{1}}u|_{\partial\overline{X}}:u\in\ker_{L^{2}}(\Delta_{g_{1}}+V_{1})\}=\{\partial_{\rho_{2}}u|_{\partial\overline{X}}:u\in\ker_{L^{2}}(\Delta_{g_{2}}+V_{2})\},$
and let $S_{i}(1)$ (associated to $\rho_{i}$) be the scattering map for the operator $\Delta_{g_{i}}+V_{i}$ for $i=1,2$. If $S_{1}(1)=S_{2}(1)$, then there exists a diffeomorphism $\psi\colon\overline{X}\to\overline{X}$ fixing the boundary pointwise, and a smooth function $c$ such that
$g_{1}=c\psi^{*}g_{2},\quad V_{1}=c^{-1}\psi^{*}V_{2}.$
The proof of [GT11]*Corollary 6.1 applies mutatis mutandis: it is not hard to see that the DN maps for the operators $\Delta_{\overline{g}_{i}}+\overline{V}_{i}$ are the same, then [CLT24]*Theorem 1.1 allows one to recover both the metrics and the potentials with the gauge mentioned in the result. The difference with [GT11]*Corollary 6.1 lies in the fact that previously only the potential could be recovered, as an application of the main theorem of [GT11].